OpenStudy (anonymous):

what are the explicit equation and domain for a geometric sequence with a first time of 4 and a second term of -12?

OpenStudy (anonymous):

@ganeshie8

OpenStudy (anonymous):

@misty1212

OpenStudy (anonymous):

helpppppp

Directrix (directrix):

4, -12,

OpenStudy (anonymous):

what's do I plug in for r??

Directrix (directrix):

-12/4 = -3 a_n = 4 * (-3) ^ (n - 1) Start with n = 1 and see if the formula yields 4. Then, let n = 2 and see if the formula yields -12

OpenStudy (anonymous):

I'm so confused

OpenStudy (anonymous):

@directrix

Directrix (directrix):

Maybe you should throw up those options for this problem so that we could decide which one it is.

OpenStudy (anonymous):

an=4(-3)^n-1; all integers where n >_ 1 an=4(-3)^n-1; all interferes where n >_ 0 an=4(36)^n-1; all integers where n >_ 1 an=4(36)^n-1; all integers where n >_ 0

Directrix (directrix):

We already know that the answer has to be A or B from the nth term formula.

Directrix (directrix):

If you put 0 in for n, you do not get the first term. So, that leaves which option as the answer?

OpenStudy (anonymous):

a??

Directrix (directrix):

Correct.